In my opinion, we ought to stop making our own drums and accept that outside suppliers offer, said Wim Niewindt, managing
Answer #1
Per Unit $      50,000 $      75,000 $    105,000
Manufactured
Direct Material 11.2 $    560,000 $    840,000 $ 1,176,000
Direct Labor (7*80%) 5.6 $    280,000 $    420,000 $    588,000
Variable Overhead (1.75*80%) 1.4 $      70,000 $    105,000 $    147,000
Fixed Company Overhead $    150,000 $    150,000 $    150,000
Rent $    157,500 $    157,500 $    157,500
Supervision Cost $      52,500 $      52,500 $      52,500
Total Manufacturing Cost $ 1,270,000 $ 1,725,000 $ 2,271,000
Lees: Purchase price from Outsider $ 1,050,000 $ 1,575,000 $ 2,205,000
Advantage $    220,000 $    150,000 $      66,000
Per Unit 50000 75000 105000
Manufactured
Direct Material 11.2 =$B$3*C1 =$B$3*D1 =$B$3*E1
Direct Labor (7*80%) =7*80% =$B$4*C1 =$B$4*D1 =$B$4*E1
Variable Overhead (1.75*80%) =1.75*80% =$B$5*C1 =$B$5*D1 =$B$5*E1
Fixed Company Overhead =3*50000 =C6 =D6
Rent 157500 =C7 =D7
Supervision Cost 52500 =C8 =D8
Total Manufacturing Cost =SUM(C3:C8) =SUM(D3:D8) =SUM(E3:E8)
Lees: Purchase price from Outsider =C1*21 =D1*21 =E1*21
Advantage =C9-C10 =D9-D10 =E9-E10